Determine the frequency of the fundamental mode and the next two harmonic that can be set up on a sitar string of length 1.0 m. Take the speed of the waves of the string to be 2.8×10^3 m/s

1
Expert's answer
2022-04-04T09:04:25-0400

"f=\frac vL=2.8~kHz,"

"f_2=\frac f2=1.4~kHz,"

"f_3=\frac f3=0.9~kHz."